AN IDEAL REFRIGERATOR REMOVES HEAT AT THE RATE OF 0.10KW FROM ITS INTERIOR (+2C) AND EXHAUSTS HEAT AT 40C. HOW MUCH ELECTRICAL POWER IS USED?


1
Expert's answer
2021-04-22T19:00:18-0400
"P'=\eta P\\P'=\frac{T_H}{T_H-T_C} P\\P'=\eta P\\100=\frac{40}{40-2} P\\P=95\ W"